OpenStudy (anonymous):

Please help me to understand why the derivative of -e^-x is e^-x. help me with the steps to get the answer. I know the derivative of e^x is e^x, don't know how to solve -e^-x. thanks

OpenStudy (anonymous):

Do you know chain rule?

OpenStudy (anonymous):

yes i know

myininaya (myininaya):

\[\frac{d}{dx}(e^u)=u'e^u\] u is a function of x

OpenStudy (anonymous):

Then you have to differentiate -e^-x as a chain rule first you differentiate e^ which is the same thing then you differenctiate -x whixh is -1 This -1 cancels the minus.

OpenStudy (anonymous):

@shankee But if i differentiate e^-x, what do i get? and what about the minus sign infront? I don't understand it please

OpenStudy (anonymous):

i know that differential of e^x is e^x but what about de differential of e^-x. I don't understand what is going on. Need some explanation.thanks

OpenStudy (anonymous):

err if you want to understand better first take e^-x and differentaiate to get the answer then just mutilpy by -1. First chain rule if you have f(g(x)) derivative= f'(g(x))*g'(x) here f(x)=e^x g(x)=-x. now differentiate using chain rule f'(g(x))=e^-x, g(x)=-1 so the net derivative is -e^-x now the question ahs an extra minus sign so multiplying by -1 removes the minus from the answer and you get only e^-x

OpenStudy (anonymous):

@shankvee Thanks alot. i get it now